OpenStudy (anonymous):

how do you find the derivative of f(x)=square root of (x-1)?

OpenStudy (lgbasallote):

express it in exponential form first (x-1)^1/2 now use power rule 1/2(x-1)^-1/2 d/dx = 1/2sqrt(x-1)

OpenStudy (anonymous):

you should perhaps memorize the derivative of \[\sqrt{x}\] because it comes up all the time. it is \[\frac{1}{2\sqrt{x}}\] you can get it using the "power rule" but it is always the same so easier just to know it. kind of like knowing that 8 times 7 is 56

OpenStudy (anonymous):

so when you take out the square root to 1/2 you don't multiply that with the x and -1 inside the parenthesis? only the outside of them, oh i see

OpenStudy (anonymous):

so \[\frac{d}{dx}\sqrt{x-1}=\frac{1}{2\sqrt{x-1}}\] and be done

OpenStudy (anonymous):

oh okay now i can substitute x value into it to find the slope. and therefore get my equation.